Oh my goodness, I'm really struggling with this one. Can anyone help me?
Consider the following sequence:
3, 12, 48, 192...
a. Write a recursive formula for the nth term of the sequence
b. Write a general formula for the nth term of the sequence c. Calculate S10 for this sequence

Answer by stanbon(75887)

Consider the following sequence:
3, 12, 48, 192...
a. Write a recursive formula for the nth term of the sequence
Ans: a(n+1) = 4*a(n)
------------------------------
b. Write a general formula for the nth term of the sequence
a(n) = 3*4^(n-1)
--------------
c. Calculate S10 for this sequence
Geometric sequence with a(1) = 3 and r = 4
----
S(10) = a(1)*r^(n-1) = 3*4^9 = 786432
--------------
Cheers,
Stan H.
------------